# GimbalDiffusion: Gravity-Aware Camera Control for Video Generation

### v1

@@ 3.2 Creating varied camera paths from 360° videos
We further introduce variations in the camera field of view (FoV). For each sequence, we sample either 1, 2, or 3 keyframes at random timestamps, with an FoV ranging from 35° to 100°. The FoV values are then interpolated using cubic splines with randomized derivative boundary conditions. Examples from our sampling are shown in fig. 3. This approach provides broad coverage of the possible viewpoints, as shown in fig. 4. Notably, this sampling slightly biases viewpoints towards the poles, which we found to better support model stability and representation of more extreme viewpoints.
Gravity annotation.
To anchor the videos within a shared global reference frame, we align their rotation with gravity. We are able to compute the absolute, gravity-aligned camera pose for each frame as
$\mathbf{E}_{\text{abs},f}=\varphi(\mathbf{R}_{\text{pano},0})\mathbf{R}_{\text{pano},f}\mathbf{E}_{\text{rel},f}\;,$ (2)
where $\varphi(\cdot)$ is a function that preserves only the roll and pitch angles (removes the yaw) of the first sampled camera pose. After this adjustment, the transformed pose $\mathbf{E}_{\text{abs},f}$ has no rotation about the global up-vector axis. This function $\varphi(\cdot)$ is obtained by computing the gravity direction in camera coordinates and inverting it to recover the corresponding up-vector, see the supplementary for the complete definition. This process establishes a gravity-aligned coordinate system that is consistent across all videos, allowing rotations to be defined within a common global frame. Consequently, each sequence is initialized at null yaw, reducing ambiguity and easing training on this representation.
Figure 4: Comparison of parameter distribution between our sampling and a typical video dataset for camera control (RealEstate10K [RealEstate10K]). Our sampling produces videos with more rotational motion, as measured by total angular distance (the sum of absolute angular differences between consecutive camera poses), and greater diversity in Euler angles (pitch, roll, yaw). All values on the $x$-axis are expressed in degrees.
Up (+90°) Forward (0°) Down (-90°)
PreciseCam
Refer to caption Refer to caption Refer to caption
w/o null pitch
Refer to caption Refer to caption Refer to caption
Ours
Refer to caption Refer to caption Refer to caption
Prompt: A prairie with grass and flowers, and a beautiful blue sky with few clouds.
Figure 5: Entanglement between prompt and camera pitch. Without a careful captioning strategy, diffusion models may ignore the camera conditioning when the prompt semantics conflict with it. Our null-pitch conditioning mitigates this issue and preserves accurate camera control.
Algorithm 1 Random camera path sampling
@@
6:for $i=1$ to NN do
7:  $d_{t}\sim\text{Uniform}(0,1)$ $\triangleright$ Random rotation duration
8:  $t_{s}\sim\text{Uniform}(0,1-d_{t})$ $\triangleright$ Random start time
9:  $t_{e}\leftarrow t_{s}+d_{t}$ $\triangleright$ End time
10:  $\mathbf{a}\sim\mathbb{S}^{2}$ $\triangleright$ Random unit axis of rotation
11:  $\theta_{\max}\sim\text{Beta}(1.0,5.0)\cdot 720^{\circ}\cdot d_{t}$ $\triangleright$ Random angular displacement
12:  Construct cubic spline $s(t)$ on $t_{s},t_{e}]$ with random boundary derivatives
13:  for $f=0$ to F1F-1 do $\triangleright$ Loop over the $F$ frames
14:   $\theta_{f}\leftarrow\text{clamp}(s(f/F),0,1)\,\theta_{\max}$ $\triangleright$ Compute the rotation angle
15:   $\mathbf{R}^{i}_{f}\leftarrow\exp([\theta_{f}\,\mathbf{a}]_{\times})$ $\triangleright$ Compute the rotation matrix
16:  end for
17:end for
18:for $f=0$ to F1F-1 do
19:  $\mathbf{R}_{\text{pano},f}\leftarrow\Pi_{i=1}^{N}\mathbf{R}^{i}_{f}$ $\triangleright$ Combine the NN rotations together
20:end for
21:return $\mathbf{R}_{\text{pano}}$
Prompt AC3D [bahmani2025ac3d] + cam. text. PreciseCam [bernal2025precisecam] + WAN  [wan2025] Ours
A luxurious cliffside resort overlooks a lush landscape, with vibrant greenery and a distant town visible beneath a bright, clear sky, evoking a sense of serene exclusivity. Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
A quiet stone church with red pews, arched windows, and historical architecture exudes reverence and solemnity through its timeless design and natural lighting. Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Prompt AC3D [bahmani2025ac3d] + cam. text. PreciseCam [bernal2025precisecam] + WAN [wan2025] Ours
A poised wingsuit flyer stands on a grassy mountain peak, overlooking a forested valley and a distant town beneath an overcast sky, evoking tension and grandeur. Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
A lively pedestrian street lined with traditional shops and trees, bathed in soft afternoon light, exudes a relaxed, vibrant energy amid a crowd of people. Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 6: Qualitative results on the SpatialVID-extreme benchmark. The input absolute camera angle is shown as a dark overlay on the image. We observe that our method aligns well with the overlay, showing the sky when the zenith (“UP”) is in the view and showing the horizon aligned with the equator line (going through “FRONT”), for example. Please refer to the supp. mat. for additional results.
3.3 Null-pitch conditioning
We observed that captioning videos using their actual rotated crops causes the model to overemphasize the textual prompt when determining the camera angle, rather than strictly adhering to the Plücker conditioning. For instance, if a prompt includes ground-level features like grass or asphalt but the camera condition is oriented toward the sky, the model tends to ignore the intended camera control and produce a downward-looking viewpoint. To mitigate this entanglement, we caption each video using an upright, forward-looking image with a 90° field of view, which strengthens the model’s adherence to the camera conditioning (see fig. 5). We refer to this strategy as “null-pitch conditioning”. Specifically, we generate two sets of crops: one using the original randomly sampled camera trajectory, and another with pitch and roll set to 0° and yaw unchanged. As illustrated in fig. 2, a trajectory that consistently points towards the ground would otherwise yield captions describing only ground features. A model trained on such captions would then expect these elements to appear in its outputs, even when the camera conditioning specifies a view of the sky, leading to semantically incorrect generations. Training with null-pitch captions instead provides descriptions of the full scene, making it easier for the model to handle extreme camera angles. During training, the randomly sampled perspective crops are used for the diffusion loss, while the null-pitch crops are used for captioning. We found that applying this scheme to 50% of the samples provides a good balance between camera controllability and prompt adherence, as shown in fig. 5.
4 Experiments
This section presents the training data, provides details on model training and losses, introduces our novel evaluation benchmark, presents baselines, and shows quantitative and qualitative experimental results on our benchmark.
4.1 Training data
We use the WEB360 dataset [wang2024360dvd], which contains geometrically calibrated panoramic videos with leveled horizons. From each 360° frame, we extract six perspective crops and apply ViPE [huang2025vipe] to obtain camera pose annotations, as detailed in Sec. 3.2. The estimated camera poses are then combined with the random crop sampling to obtain the final camera poses, as described in Sec. 3.2.
During training, we sample a camera trajectory using Algorithm 1. We use InternVL-3-2B [zhu2025internvl3] for captioning, providing it with 6 evenly spaced frames across the video timeline. This is also done during training since new videos from different trajectories are sampled each iteration.
Table 1: Quantitative results on our SpatialVID-extreme evaluation dataset. Our method outperforms all baselines on both relative and absolute camera rotation metrics while achieving competitive performance on content alignment (CLIP) and aesthetics (FID, FVD).
Method PitchErr. (abs.) $\downarrow$ GravityErr. (abs.) $\downarrow$ RotErr (rel.) $\downarrow$ TransErr $\downarrow$ CLIP $\uparrow$ FID $\downarrow$ FVD $\downarrow$
@@ 4.1 Training data
[…]
4.2 Model and losses
We adopt the same lightweight ControlNet architecture as AC3D [bahmani2025ac3d], using CogVideoX-2B [yang2024cogvideox] as the backbone diffusion model. Following their method, we train the model only on the first 40% of diffusion timesteps and use 8 layers of the original network for the ControlNet training, leaving the remaining layers frozen. We use the same training loss with normalization [karras2024analyzing], and change the relative Plücker-ray representation to our absolute Plücker representation for camera control. We train on four NVIDIA A100 GPUs with 80GB VRAM until convergence ($\sim 8,000$ iterations). See the supplementary material for details and results with our method trained on the WAN [wan2025] backbone.
4.3 Evaluation benchmark
There exists no standard evaluation dataset containing paired text prompts, camera poses, and reference videos that provide extreme camera angles. Most methods [bahmani2025ac3d, he2025cameractrl] use test splits from video datasets like RealEstate10K [RealEstate10K], which have very limited diversity in terms of pitch and roll angles, as shown in fig. 4 and the supplementary material.
We propose a new evaluation benchmark, featuring diverse, high-quality videos, distributed with near-uniform pitch and roll coverage. We start with 371K video clips from the “HQ” subset of the SpatialVID [wang2025spatialvidlargescalevideodataset] dataset, which still exhibit a bias toward forward-looking orientations. We estimate the average pitch from the videos using Perspective Fields [jin2023perspective] and organize them into 10-degree bins spanning -85 to +85 degrees. We manually filter out videos where the pitch estimation clearly failed until we obtain a subset of 140 videos with a uniform pitch distribution. For the roll, we found that even among the 371K videos, nearly all have near-zero roll. We thus artificially add a roll angle by randomly sampling a roll trajectory between -40° and 40° and applying it by warping the video. We name the resulting dataset SpatialVID-extreme. By having a diverse set of pitches and rolls, we ensure a diverse set of text prompts and reference videos for FID and FVD computations. Please refer to the supplementary material for an evaluation of the diversity of our benchmark dataset.
For the camera trajectories fed to the models, we reuse the original camera pose translations, but discard the rotation. We instead generate a random rotation by sampling roll, pitch, and yaw trajectories (see the supp. for details). This strategy allows for a broader range of rotations and also better tests the methods’ adherence to the camera trajectory. Indeed, we receive a normal environment description as input prompt, but for extreme camera angles, such as looking straight up, adhering to the camera conditioning likely requires generating views of the sky.
Table 2: Ablation study on null pitch conditioning. Incorporating null-pitch conditioning improves performance across all metrics and enhances the model’s robustness.
Method PitchErr. (abs.) $\downarrow$ GravityErr. (abs.) $\downarrow$ RotErr (rel.) $\downarrow$ TransErr $\downarrow$ CLIP $\uparrow$ FID $\downarrow$ FVD $\downarrow$
w/o null pitch conditioning 28.75 31.70 21.55 0.76 21.29 113.90 983.50
[…]
4.4 Baselines and metrics
PreciseCam + WAN.
A straightforward approach to camera-controlled video generation is to first use a camera-conditioned image generator to synthesize the initial frame, and then apply an image-to-video model to produce the full sequence. This baseline follows that strategy: we use PreciseCam [bernal2025precisecam] to generate the first frame conditioned on the absolute pitch, roll, and field of view, and then use a relative camera control model, WAN-2.1-Fun-I2V Camera Control [wan2025], to extend it into a video.
AC3D + cam. text.
We employ AC3D [bahmani2025ac3d], a text-to-video model with camera control, as baseline. We use their pre-trained checkpoint that uses the same CogVideoX-2B as our model. To provide it with the required initial and final camera poses, we append a text description of the absolute poses, generated from a fixed template, to the original content prompt (refer to the supp. for details).
AC3D + cam. text. + abs. Plücker
Metrics.
To evaluate camera rotation controls, we first measure absolute orientation accuracy using two metrics: angular pitch error, and gravity error, defined as the angle between the ground-truth up vector and the up vector of the generated video, in camera space. For each frame, we extract the pitch and up vector using Perspective Fields [jin2023perspective]. We do not directly evaluate roll error, as estimating it near the poles (looking straight up or down) is unreliable; however, roll deviations are implicitly reflected in the gravity error.
To evaluate the relative rotation error (RotErr), we run VGGT on the generated frames and compute the average relative angular error. We reuse the same definition of rotation error as [he2024cameractrl]. We evaluate content adherence using the CLIP score [hessel2021clipscore] to measure the alignment between the generated video frames and the reference text description. We employ the Fréchet Inception Distance (FID) and Fréchet Video Distance (FVD) to assess the visual quality and realism of the generated frames and sequences, respectively. To evaluate camera motion accuracy, we compute a translation error by comparing the estimated and ground-truth trajectories; both trajectories are normalized prior to computing the error to account for scale differences.
@@ 4.5 Results
We present the result of our evaluation in Table 1, which shows that our method clearly outperforms existing methods. Without any change to the architecture, our method nearly halves the pitch and gravity errors and significantly reduces the relative rotation error. We attribute the slight decrease in performance on CLIP, FID and FVD scores to differences in the training data: AC3D is trained on all 65k videos and prompts from RealEstate10K, whereas we only train on 2K 360 videos from WEB360.
We showcase qualitative results in Fig. 6. We observe that while AC3D follows the motion correctly (e.g., the camera progressively looks down throughout the video), it fails to capture the initial target view direction. Feeding it instead the absolute camera poses, which is out of its training distribution, results in negligible improvements in absolute metrics but generates noticeable artefacts (see supp.). PreciseCam [bernal2025precisecam] coupled with WAN-I2V [wan2025] exhibits several failure cases, in particular when used with extreme or unusual camera angles. This is because the input prompt describes the overall environment, which may not be what the camera is pointed at. For example, the prompt in fig. 6 (top left) describes ground elements while the camera aims at the sky. We observe that PreciseCam focuses on text content and often fails to produce the correct viewpoint when it can’t place the content in a semantically meaningful way within the image. In contrast, we use our null-pitch conditioning to ensure that generic descriptions of environments do not prevent the generation of accurate camera viewpoints, as shown in fig. 6 (right).
@@
We ablate our proposed null-pitch conditioning in Tab. 2, which shows that removing it leads to a 20% and 17% drop in absolute camera pitch and gravity accuracy, respectively. This is corroborated qualitatively in fig. 5: in the top and middle rows, where null-pitch conditioning is omitted, the model incorrectly introduces elements such as grass even when the camera is oriented fully upward.

### v2

@@ 3.2 Creating varied camera paths from 360° videos
We further introduce variations in the camera field of view (FoV). For each sequence, we sample either 1, 2, or 3 keyframes at random timestamps, with an FoV ranging from 35° to 100°. The FoV values are then interpolated using cubic splines with randomized derivative boundary conditions. This approach provides much broader coverage of possible viewpoints than existing datasets (see supplementary). Notably, this sampling slightly biases viewpoints towards the poles, which we found to better support model stability and representation of more extreme viewpoints. Examples from our sampling are shown in fig. 3.
Algorithm 1 Random camera rotation path sampling
$\text{pitch}_{0}\sim\text{Uniform}(-90\degree{},90\degree{})$ ; $\text{roll}_{0}\sim\text{Uniform}(-90\degree{},90\degree{})$ ; $\text{yaw}_{0}\sim\text{Uniform}(0\degree{},360\degree{})$
$\mathbf{R}_{\text{pano},f}\leftarrow\left(\mathcal{R}_{\text{YXZ}}(\text{yaw}_{0},\text{pitch}_{0},\text{roll}_{0})\right)_{f=0}^{F-1}$ $\triangleright$ Initialize the camera orientation
$N\sim\{0,1,2,3\}$ $\triangleright$ Random number of rotation axis
for $i=1$ to NN do
  $d_{t}\sim\text{Uniform}(0,1)$ $\triangleright$ Random rotation duration
  $t_{s}\sim\text{Uniform}(0,1-d_{t})$ ; $t_{e}\leftarrow t_{s}+d_{t}$ $\triangleright$ Random start and end time
  $\mathbf{a}\sim\mathbb{S}^{2}$ $\triangleright$ Random unit axis of rotation
  $\theta_{\max}\sim\text{Beta}(1.0,5.0)\cdot 720^{\circ}\cdot d_{t}$ $\triangleright$ Random angular displacement
  Construct cubic spline $s(t)$ on $t_{s},t_{e}]$ with random boundary derivatives
  for $f=0$ to F1F-1 do $\triangleright$ Loop over the $F$ frames
   $\theta_{f}\leftarrow\text{clamp}(s(f/F),0,1)\,\theta_{\max}$ $\triangleright$ Compute the rotation angle
   $\mathbf{R}_{\text{pano},f}\leftarrow\exp([\theta_{f}\,\mathbf{a}]_{\times})\times\mathbf{R}_{\text{pano},f}$ $\triangleright$ Update the camera orientation
  end for
end for
return $\mathbf{R}_{\text{pano}}$
UCPE PreciseCam+GEN3C AC3D+cam. text. Ours UCPE PreciseCam+GEN3C AC3D+cam. text. Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
A sleek, modern interior features a dark, hexagonally lit ceiling that radiates brightness and sophistication, creating a visually striking and inviting atmosphere. A modern skyscraper rises against a clear blue sky, its grid-like facade reflecting sunlight in a bustling urban setting.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
A serene forest is bisected by a powerful waterfall, its vivid greens, yellows, and reds evoking tranquility and natural majesty. A sleek, modern home with high ceilings, natural light, and contemporary furnishings features a man in a blue suit in its open-plan living area.
Figure 4: Qualitative results on the SpatialVID-extreme benchmark. The input absolute camera angle is shown as a dark overlay on the image. We observe that our method aligns well with the overlay, showing the sky when the zenith (“UP”) is in the view and showing the horizon aligned with the equator line (going through “FRONT”), for example. Please refer to the supp. for additional results.
3.3 Null-pitch conditioning
We observed that captioning videos using their actual rotated crops causes the model to overemphasize the textual prompt when inferring the camera angles, rather than strictly adhering to the Plücker conditioning. This entanglement hinders the achievement of extreme camera angles, where visual content diverges most from the prompt. For instance, if a prompt describes ground features such as grass but the camera should face the sky, the model often still outputs a downward-looking view. To mitigate this entanglement, we caption each video using an upright, forward-looking image with a 90° field of view, thereby strengthening the model’s adherence to the camera conditioning. We refer to this strategy as “null-pitch conditioning”. Specifically, we generate two sets of crops (see fig. 2): one using the original randomly sampled camera trajectory, and another with pitch and roll set to 0° (with yaw unchanged). As illustrated in fig. 2, a trajectory that consistently points towards the ground would otherwise yield captions describing only ground features. A model trained on such captions would then expect these elements to appear in its outputs, even when the camera conditioning specifies a view of the sky, leading to semantically incorrect generations. Training with null-pitch captions, which provide descriptions of the full scene, makes it easier for the model to handle extreme camera angles. During training, the randomly sampled perspective crops are used for the diffusion loss, while the null-pitch crops are used for captioning.
While this strategy helps the video model in handling extreme angles, we found it still struggled with undesirable elements present directly below the camera in real 360° training videos, such as tripods or the hand of a person holding a selfie stick. Indeed, since these elements are never present in the training caption, the model learns to generate these objects as if they were normal scene content. To prevent this, we generate a third set of videos (not shown in fig. 2), this time looking straight down (-90° pitch), and caption them with the VLM. We append the resulting “look-down” caption to the null-pitch caption 50% of the time. At inference time, this “look-down” prompt is used in the classifier-free guidance as a negative prompt to ensure that undesirable content is absent from the generated videos. Please see the supplementary for additional details.
4 Experiments
This section presents the training data, details model training and losses, introduces our two novel evaluation benchmarks, presents baselines, and shows quantitative and qualitative experimental results on our benchmarks.
4.1 Training data
We use the “YouTube videos” subset of the PanoVid dataset [43], which contains a diverse set of high-quality 360° videos, and filter out videos where the shortest side is less than 900 pixels. We obtain relative camera poses and per-frame up-vectors using the procedure from Sec. 3.2. We discard videos where pose estimation either failed or resulted in extreme mean acceleration values, typically indicating unreliable camera poses. Our filtering strategy reduces the number of training videos from 7797 to 4798 with high-quality camera poses (38% discarded).
During training, we sample a random camera trajectory using Algorithm 1 at a resolution of $720\times 480$, effectively acting as an infinite data augmentation strategy. We use InternVL-3-2B [52] for captioning, providing it with 6 evenly spaced frames across the video timeline. This is also done during training since new videos from different trajectories are sampled each iteration. Prompt lengths are further sampled between short, medium and long.
4.2 Model and losses
We start from the WAN-2.2 5B text-to-video model [34], and add a lightweight camera encoder consisting of two convolution layers and a group normalization layer (see supp. for details). The last convolution is zero-initialized to avoid instabilities at the beginning of training. Its output is then added after initial patchify layer of the DiT backbone.
During training, both the camera encoder and the backbone are finetuned for 30k iterations (1 day on 8 NVIDIA H200 GPUs). We use the AdamW optimizer with a learning rate of $10^{-5}$ and a weight decay of 0.01.
4.3 Video camera control benchmark
There exists no standard evaluation dataset containing paired text prompts, camera poses, and reference videos that provide extreme camera angles. Most methods [2, 14] use test splits from video datasets like RealEstate10K [10], which have very limited diversity in terms of pitch and roll angles, as shown in the supplementary material.
We propose a new evaluation benchmark, featuring diverse, high-quality videos, distributed with near-uniform pitch and roll coverage. We start with 371K video clips from the “HQ” subset of the SpatialVID [35] dataset, which still exhibit a bias toward forward-looking orientations. We estimate the average pitch from the videos using Perspective Fields [19] and organize them into 10-degree bins spanning -85° to +85°. We manually filter out videos where the pitch estimation clearly failed until we obtain a subset of 140 videos with a uniform pitch distribution. For the roll, we found that even among the 371K videos, nearly all have near-zero roll. We thus artificially add a roll angle by randomly sampling a roll trajectory between -40° and 40° and applying it by warping the video. We name the resulting dataset SpatialVID-extreme. By having a diverse set of pitches and rolls, we ensure a diverse set of text prompts and reference videos for FID and FVD computations. Please refer to the supplementary material for an evaluation of the diversity of our benchmark dataset.
For the camera trajectories fed to the models, we annotate translations in metric scale using ViPE [17], but discard the rotation. We instead generate a random rotation by sampling roll, pitch, and yaw trajectories (see the supp. for details). This strategy allows for a broader range of rotations and also better tests the methods’ adherence to the camera trajectory. Indeed, we receive a normal environment description as input prompt, but for extreme camera angles, such as looking straight up, adhering to the camera conditioning likely requires generating views of the sky.
4.4 Baselines and metrics
UCPE
[50] is a concurrent method that also generates videos in a gravity-aligned reference frame, using WAN-2.1 as the backbone. Unlike ours, UCPE requires a limited auxiliary dataset to simulate rotation diversity within panoramas, resulting in a less diverse dataset than our proposed approach.
AC3D+cam. text.
We employ AC3D [2], a text-to-video model with camera control, as baseline. We use their pre-trained checkpoint using CogVideoX-2B as the base video model. To provide it with the required initial and final camera poses, we append a text description of the absolute poses, generated from a fixed template, to the original content prompt (refer to the supp. for details).
AC3D+cam. text.+abs. Plücker.
PreciseCam+WAN.
A straightforward approach to camera-controlled video generation is to first use a camera-conditioned image generator to synthesize the initial frame, and then apply an image-to-video model to produce the full sequence. This baseline follows that strategy: PreciseCam [5] generates the first frame conditioned on the absolute pitch, roll, and field of view, and then WAN-2.1-Fun-I2V Camera Control [34] prolongs that first frame to a video.
PreciseCam+GEN3C.
GEN3C [30] takes as input a 3D point cloud from a single image or an input video which can then be rendered from novel camera views. We reuse the original implementation which predicts the depth using MoGE [38] and feed it the first frame from PreciseCam. We then project the point cloud along the relative camera poses and feed them to GEN3C. Note however that GEN3C can only generate novel views of static scenes, and therefore cannot produce dynamic video content (e.g., walking people).
Metrics.
To evaluate camera rotation controls, we first measure absolute orientation accuracy using two metrics: angular pitch error (“PitchErr”), and gravity error (“GravityErr”), defined as the angle between the ground-truth up vector and that of the generated video, in camera space. The pitch angle and up vector are obtained from each frame using Perspective Fields [19]. Roll error is not directly computed, as estimating it near the poles (looking straight up or down) is unreliable; however, roll deviations are implicitly reflected in the gravity error.
To evaluate the relative rotation error (“RotErr”), we run VGGT [36] on the generated frames and compute the average relative angular error. We reuse the same definition of rotation error as [13]. To evaluate camera motion accuracy, we compute a translation error (“TransErr”) by comparing the estimated and ground-truth trajectories; both trajectories are normalized prior to computing the error to account for scale differences. The CLIP score [15] is computed to measure the alignment between the generated video frames and the reference text description. Finally, we employ the Fréchet Inception Distance (FID) and Fréchet Video Distance (FVD) to assess the visual quality and realism of the generated frames and sequences, respectively.
Table 1: Quantitative results on our SpatialVID-extreme evaluation dataset. Our method outperforms all baselines on absolute camera orientation metrics (PitchErr and GravityErr), and all dynamic video generation methods on relative rotation error (RotErr), while achieving competitive performance on translation error (TransErr), and aesthetics (FID, FVD). Furthermore, removing the null-pitch conditioning drastically degrades absolute orientation metrics.
Method PitchErr GravityErr RotErr TransErr CLIP FID FVD
Novel view synthesis methods
PreciseCam+GEN3C 26.82 31.09 10.99 0.40 20.17 101.41 830.7
Dynamic video generation methods
@@ 4.5 Results
We present the quantitative results of our evaluation in Sec. 4.4. Overall, our method outperforms all dynamic video generation baselines on all camera orientation metrics (“PitchErr”, “GravityErr”, and “RotErr”). Since GEN3C can only synthesize novel viewpoints without scene motion, we include its results here for reference only. Our method achieves FID scores similar to those of other baselines and FVD scores on par with UCPE and AC3D. Baselines using PreciseCam as a starting point improve the FVD score, at the cost of much worse camera orientation metrics. We observe that our method results in a slight drop in CLIP score, see the discussion on prompt-camera entanglement in Sec. 4.6 below.
Representative qualitative results are illustrated in fig. 4 (see the supplementary for more). We observe that all baselines struggle with extreme angles (e.g., the “up” starting point in the bottom-left example). Here, they attempt to generate visual content in accordance with the prompt, but this contradicts the desired camera orientation. Another example is shown in the top-right example, where the “skyscraper rising against a clear blue sky” eventually becomes visible in “Ours” (as it should when the camera starts to look up), whereas both UCPE and “AC3D+cam. text” are incorrectly rendering it right from the first frame. In this example, PreciseCam fails to properly generate the first frame. Overall, visual examples confirm that our approach provides the best camera control, even in the case of extreme viewpoints.
@@
We ablate our proposed null-pitch conditioning in Sec. 4.4, which shows that its absence leads to a 67% and 52% increase in camera pitch and gravity error, respectively. Next, we investigate the effect of null-pitch conditioning.
4.6 Prompt-camera entanglement evaluation
To investigate the effect of null-pitch conditioning, we develop a second benchmark dataset to evaluate whether methods can disambiguate between two possibly conflicting inputs: the prompt and the camera pitch angle. Indeed, pitch is highly correlated with semantic concepts, e.g., a +90° pitch (looking up) is strongly associated with images of the sky/ceiling, whereas a -90° pitch is strongly associated with ground features. Here, roll and yaw angles are ignored, since both are uncorrelated to object semantics in the field of view.
We start by randomly selecting 20 panoramas from the PolyHaven dataset [12]. For each, we take a forward-looking crop with a 90° FoV and caption it using InternVL-3-8B [52]. We then give this prompt to each evaluated method along with a static camera orientation. The methods are evaluated on pitch angles going from -90° to +90°, in 10° increments. In total, this benchmark consists of 380 pairs of input prompts and pitch angles.
The results obtained with baselines providing gravity-aligned camera control (UCPE [50] and PreciseCam [5]) are shown in fig. 5. Qualitatively, we observe in fig. 5-(a) that our proposed null-pitch conditioning enforces strict adherence to the camera orientation when the camera is pointing straight up or straight down. Quantitatively, we observe in fig. 5-(b) (top) a substantially lower pitch-angle error across all 20 scenes, particularly at extreme absolute pitch values.
One simple way to assess the visual alignment between the text prompt and the output video is to compute the CLIP score directly between them. However, this naive approach has a key limitation: the prompt features may rightly be absent from the video if they violate the user-specified camera constraint (e.g., textual descriptions of ground content should not be visible if the camera is pointing up)—but not doing so would negatively affect the CLIP score. To address this contradiction, we compute the CLIP score between the generated video frames and a caption of the ground-truth frame (obtained with InternVL-3-8B) for each pitch angle separately, and report results in fig. 5-(b) (bottom). We observe that null-pitch conditioning improves CLIP similarity at extreme pitch values, aligning results closer to the ground-truth curve. See the supplementary material for a detailed analysis.
PreciseCam UCPE w/o null pitch Ours Ground truth
Up (+90°)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Forward (0°)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Down (-90°)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A serene landscape unfolds with dry grass swaying gently in the foreground. The sun shines brightly in the clear blue sky, casting a warm glow over the scene. Sparse green bushes dot the landscape, with distant hills visible on the horizon. The scene remains tranquil, with no visible movement or change in the environment.”
(a) Qualitative example from our entanglement benchmark
Refer to captionRefer to caption
(b) Pitch error and CLIP variation over our entanglement benchmark.
Figure 5: Evaluating the entanglement between prompt and camera pitch. (a) Without a careful captioning strategy, generative models may ignore the camera conditioning when it conflicts with the prompt semantics, which is especially problematic at extreme pitch values (e.g., -90° and +90°). (b) Quantitative evaluation of this effect: our null-pitch conditioning strategy reduces the aggregated pitch error at extreme angles (top) and maximizes the content similarity with the ground truth, as reflected in the CLIP scores (bottom). Here, the CLIP scores are shifted to each method’s average to highlight their variations with respect to the input pitch. Refer to the supp. for additional results.